# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

### v2

@@ 7 Discussion
We have introduced a versatile and modular library designed to advance the application of deep learning to RNA structural analysis. By providing a suite of diverse and well-defined datasets and splits, our library enables robust benchmarking and facilitates the development of innovative computational models. It ensures reproducibility and promotes standardized evaluation, fostering confidence in computational findings. Its flexibility allows researchers to easily build additional tasks specific to their research needs in a wide range of RNA structural studies.
In the future, the development of additional tasks, such as the assessment of structural models (Townshend et al., 2021c) and a greater focus on RNA embeddings, presents exciting opportunities. As the understanding of the dynamic nature of biological macromolecules evolves, so may the preferred representation for RNA. Our tool’s built-in 2.5D-graph representations offer extensive adaptability and have been used to develop recent competitive RNA structure-based models (Wang et al., 2024). Nevertheless, the library can easily be extended to incorporate new or improved representations as they emerge.
In this work, we unify various biological challenges where RNA structure is a key player, and provide direct access to high quality data and benchmarking functionality for deep learning model development and computational biology communities alike. Furthermore, the choice of integrating the benchmark in the rnaglib toolkit opens access to non-specialist researchers, and provides robust guarantees for quality and reproducibility. We hope that this benchmark will facilitate the development of new deep learning tools in the highly promising area of RNA structural biology.
Appendix A Technical Appendices and Supplementary Material
A.1 Details about datasets
In this appendix, we first detail more extensively the datasets and data preprocessing applied (Section A.1). Then, we provide additional results of the models tested on our tasks together with details about training and hyperparameters used to perform the experiments (Section A.2).
A.1.1 Datasets and general preprocessing
@@
As mentioned in section 3, RNA molecules present in the PDB files exhibit a bimodal distribution over the number of residues. Figure 5(a) displays the distribution of the RNA sizes (defined as the number of residues) of the RNAs from the PDB.
@@
A single system, such as rRNA, can gather several independent, non-interacting RNA chains in a single PDB file. In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components. Figure 5(c) represents the distribution of the number of residues by connected components.
We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of the data loading as well as model training. To this end, we filter RNA chunks and only keep the ones having between 15 and 500 nucleotides. Figure 5(c) displays the distribution of RNA chunks sizes after filtering.
The redundancy removal process we apply is the following. First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (relying on a sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program). Then, within each cluster, we select the sample with the highest resolution. Then we perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure. Afterwards, when instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering"). Since we only select one representative sample per cluster, the number of clusters prefigures the number of samples of the final dataset. Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
For this reason, we study the impact of both the redundancy removal and the splitting clustering threshold on the number of clusters generated in the case of RNA-CM task. Results are reported in figure 6. Figure 6(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1. The chosen value is 0.90. Figure 6(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold. In all the experiments, the splitting clustering is performed based on US-align similarity (structural similarity).
@@
A.1.2 Task-specific data preprocessing
@@
RNA-Ligand
We selected the three most frequent ligands which were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022]. We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework. The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamicin C1A (LLL) and aminoglycoside TC007 (8UZ). Their structures are displayed in Figure 7.
(a) Paromycin (PAR)
(b) Gentamycin C1A (LLL)
@@
Figure 7: Structures of the ligands selected for RNA-Ligand task
RNA-GO
@@
When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to "structural constituent of ribosome", "ribosome" and "tRNA"). We also remove the GO-terms which are underrepresented (those which occur less than 50 times in our experiments). We then remove the GO-terms which are very correlated to other ones by performing GO-terms clustering based on correlation matrix (with correlation threshold 0.9) and keeping only one representative GO-term per cluster. After this preprocessing, 5 GO-terms are remaining: 0000353 (formation of quadruple SL/U4/U5/U6 snRNP), 0005682 (U5 snRNP), 0005686 (U2 snRNP), 0005688 (U6 snRNP) and 0010468 (regulation of gene expression).
A.2 Additional Results

### v3

@@ 7 Discussion
We have introduced a versatile and modular library designed to advance deep learning for RNA structural analysis. By standardizing datasets, splits, and evaluation as benchmarking tasks, the library enables reproducible comparisons and lowers the barrier for developing new computational models. Its standardized workflows foster confidence in results, and its modularity makes it straightforward for researchers to extend the suite with additional tasks.
Our experiments show that RNA tasks are both learnable and challenging, and that representation choice matters: coarse-grained 2.5D graphs provide strong baselines, while 3D encoders capture additional signal beyond sequence-only models. These findings offer practical guidance for model development and underscore the importance of well-founded splitting strategies.
Looking forward, the benchmark can be extended to new directions such as structural model assessment (Townshend et al., 2021c), transferable RNA embeddings, and the integration of dynamic conformational ensembles. The built-in 2.5D representations have already enabled competitive RNA models (Wang et al., 2024) and hold further promise, while the modular design ensures that new and improved representations can be seamlessly integrated as they emerge.
By consolidating diverse RNA structure–function challenges into a single, extensible library, we create an accessible common foundation for future research in the field. We hope that this benchmark will facilitate the development of new deep learning tools in the highly promising area of RNA structural biology.
Reproducibility Statement
All results can be reproduced with the code provided in the experiments repository. All plotting scripts are also available, along with precomputed results files. We also provide two simple scripts to retrain all models.
The code to use the benchmark and the proposed tasks is extensively documented. A multitude of tutorials and how-tos exist with regards to all parts of the package. The tool can be easily installed as a python package and is maintained on Github. Additional and complete information on preprocessing is available in Appendix  B. Results can be reproduced with limited compute as specified in Appendix  D.
Funding
This project was provided with computing and storage resources by GENCI at IDRIS thanks to the grant 2025-AD010315435R1 on the supercomputer Jean Zay’s V100/A100/ H100 partition.
Appendix
In this Appendix, we first clearly differentiate the contribution of the current work with the previously published rnaglib in Section A. We then provide analysis of the initial data, a justification of the generic data preprocessing, and details of extra-steps taken for specific tasks (Section B). In Section C, we provide a data analysis of the resulting datasets obtained for each of our tasks. In Section D, we analyze computational requirements of training models on our benchmarks. In Section F, we report additional details about the representation and models used in the paper, as well as about the training parameters used for the models presented in the main body of the paper. Finally, we provide additional results of the models tested on our tasks (Section G).
Appendix A Positioning with regards to rnaglib
The proposed tool is integrated within a pre-existing rnaglib python package [Mallet et al., 2022]. In this Section we expand on the differences between the two.
The pre-existing package consists of a database of pre-computed 2.5D graphs, lowering the barrier to entry for people interested in using this representation in their applications. In addition, it proposed a set of methods to compare 2.5D graphs using a customized graph-edit distance, plotting scripts, and a kernel to compare 2.5D rooted subgraphs. Finally, it proposed a machine-learning model compatible with this representation, with metric-learning pretraining on the kernels, as proposed in RNAmigos Oliver et al. [2020].
Our current submission significantly expands on these fundamentals. Notably, we introduce all the infrastructure necessary to build a task. We introduce feature computers to easily integrate additional features and annotations, such as RNA language model embeddings. We create a representation abstraction and collating routing to go beyond the exclusive use of 2.5D representation. We also add a dataset object that can be saved or downloaded from Zenodo, as well as routines to remove redundancy and split a dataset. All of these abstract objects are designed to be flexible and modular so that implementing a representation or splitting strategy smoothly fits in our pipeline.
We then implement concrete tools using this infrastructure. Examples include GO term annotators, USAlign splitters or support for sequence, 3D point clouds, voxel grids or 3D graphs representations, and are detailed in Section 3. Finally, we propose our benchmark tasks taking advantage of those tools (see Section 4), as well as results leveraging this benchmark (see Section 6).
Appendix B Datasets Processing
B.1 General preprocessing
@@
As mentioned in Section 3, RNA molecules in the PDB files exhibit a bimodal distribution over the number of residues. Figure 5(a) displays the distribution of the RNA sizes (defined as the number of residues) of the RNAs from the PDB.
@@
A single system, such as rRNA, can include several independent, non-interacting RNA chains in a single PDB file. In order to work at a more biologically relevant scale, we partition the raw RNAs from PDB files into connected components. Figure 5(c) shows the distribution of the number of residues by connected components.
We remove the RNA components of insufficient size for structure-based machine learning, as well as very large components which negatively impact the computational performance of data loading as well as model training. To this end, we filter RNA chunks and only keep those with between 15 and 500 nucleotides. Figure 5(c) displays the distribution of RNA chunks sizes after filtering.
Our redundancy-removal process is as follows. First, we perform sequence-based clustering based on a similarity threshold above which RNA fragments are clustered together (using the sequence similarity metric computed using CD-HIT [Fu et al., 2012] clustering program). Then, within each cluster, we select the sample with the highest resolution. We then perform structure-based clustering and structure-based redundancy removal (relying on the TM-Score, a structural similarity metric, computed using US-align platform [Zhang et al., 2022a]) following the same procedure. When instantiating the splitters, we perform structure-based clustering with a different threshold to define the clusters which will be required to be grouped either in train, val or test set (which we call "splitting clustering"). Since we only select one representative sample per cluster, the number of clusters determines the number of samples of the final dataset. Therefore, a tradeoff is to be considered between having a large amount of data and discarding redundant RNAs.
For this reason, we study the impact of both redundancy removal and splitting clustering threshold on the number of clusters generated in the case of the RNA-CM task. Results are reported in Figure 6. Figure 6(a) displays, for 4 different sequence-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. Here, the redundancy removal thresholds 0.90, 0.80 and 0.70 give the same plot since CD-HIT similarity values are strongly concentrated around 0.5 and 1. The chosen value is 0.90. Figure 6(b) displays, for 4 different structure-based redundancy removal thresholds (each with a different color), the scatter plot of the number of clusters obtained based on the structure-based splitting clustering threshold. In all cases, the structure-based redundancy removal is performed after a sequence-based redundancy removal using a 0.90 threshold. In all experiments, the splitting clustering is performed based on US-align structural similarity.
@@
Our library permits setting different structure and sequence similarity thresholds for splitting, depending on use case and data availability. In Figure 7, we plot the performance of our baseline model on the redundant version of RNA-Site, varying the similarity threshold used. Our results show that more stringent splitting leads to a decreased, more realistic performance. The trade-off between leakage and learning signal can thereby be tweaked by practitioners. We point out that structure and sequence thresholds are not directly comparable, since their calculation differs. Setting a threshold of 1 results in isolated points for our structure similarity metric, resulting in comparable results to random splitting. This is in contrast with our sequence-based similarity metric, whose value is exactly one for highly similar structures. Using either a non-redundant dataset or rigorous splitting, alleviates the risk of data leakage, and using both maximizes the reliability of the results.
Figure 7: In a redundant setting, stricter splitting thresholds reduce data leakage, shown for RNA-Site.
B.2 Task-specific Preprocessing
@@
RNA-Ligand We selected the three most frequent ligands that were neither modified RNA or DNA residues nor modified protein residues, following the ligand definition proposed by Hariboss [Panei et al., 2022]. We only retained the binding pockets binding to one of these three ligands in order to ensure the amount of binding pockets binding to each ligand would be significant enough to enable learning in a multi-class classification framework. The three ligands retained are paromomycin (called PAR in the PDB nomenclature), gentamicin C1A (LLL) and aminoglycoside TC007 (8UZ). Their structures are displayed in Figure 8.
(a) Paramomycin (PAR)
(b) Gentamicin C1a (LLL)
@@
Figure 8: Structures of the ligands selected for the RNA-Ligand task
@@
RNA-GO When building RNA-GO, we explore all the GO-terms of RNAs from the PDB, remove the GO-terms which occur more than 1000 times (these are in fact the GO-terms corresponding to "structural constituent of ribosome", "ribosome" and "tRNA"). We also remove the GO-terms which are underrepresented (those which occur less than 50 times in our experiments). We then remove the GO-terms which are very correlated to other ones by performing GO-terms clustering based on correlation matrix (with correlation threshold 0.9) and keeping only one representative GO-term per cluster. After this preprocessing, 5 GO-terms are remaining: 0000353 (formation of quadruple SL/U4/U5/U6 snRNP), 0005682 (U5 snRNP), 0005686 (U2 snRNP), 0005688 (U6 snRNP) and 0010468 (regulation of gene expression).
Appendix C Data Analysis
Figures 910 show the family distributions of our proposed datasets, annotated with both Rfam and NAKB labels. We have chosen Rfam clans and NAKB annotations as the right level of granularity, since they provide a good overview of the present functions, without being overly detailed or too general. Across tasks we observe substantial functional diversity, with the exception of RNA-Ligand, which was pre-processed to include only RNAs known to bind one of three ligands. Overall, NAKB provides slightly broader coverage of annotated RNA structures compared to Rfam. Importantly, the Rfam-annotated RNAs are not a strict subset of those annotated by NAKB, so the true combined coverage is somewhat higher. To our knowledge, however, there is no standardized mapping between Rfam clans and NAKB annotations.
A large fraction of the structures correspond to ribosomal complexes, particularly ribosomal RNAs and tRNAs. This bias reflects the composition of the PDB, where many RNA crystal structures originate from RNAs crystallized in complexes with proteins. Given that the ribosome is a central molecular machine of the cell and has been extensively studied by structural biologists, its overrepresentation in our data is expected.
Table 2 reports dataset statistics, highlighting the wide range of RNA sizes present in our collection. Taken together, the diversity of RNA functions and sizes represented in our datasets ensures that models trained on them are likely to generalize across RNA types.
Table 2: We provide statistics on the size distribution of RNA samples in our datasets. Note that depending on the task, a size cutoff at either 300 or 500 nucleotides is applied.
Task Name Total Nodes Dataset Size Min Nodes Max Nodes Mean Nodes Median Nodes RNAs $\geq$200 RNAs $\geq$200 (%) Num Classes Num Features
rna_cm 11,208 197 16 482 56.89 40.00 4 2.0 2 4
rna_go 46,557 499 10 299 93.30 93.00 3 0.6 5 4
rna_if 164,769 2,729 16 298 60.38 44.00 72 2.6 5 1
rna_ligand 14,632 290 27 78 50.46 49.00 0 0.0 3 4
rna_prot 74,257 1,274 16 488 58.29 40.50 36 2.8 2 4
rna_site 14,430 224 16 478 64.42 47.00 9 4.0 2 4
Refer to captionRefer to caption
(a) CM
Refer to captionRefer to caption
(b) GO
Refer to captionRefer to caption
(c) IF
Refer to captionRefer to caption
(d) Ligand
Refer to captionRefer to caption
(e) Prot
Refer to captionRefer to caption
(f) Site
Figure 9: Distribution of NAKB functional annotations across task datasets. For each dataset, the top plot excludes unannotated RNAs and the bottom includes them.
Refer to captionRefer to caption
(a) CM
Refer to captionRefer to caption
(b) GO
Refer to captionRefer to caption
(c) IF
Refer to captionRefer to caption
(d) Ligand
Refer to captionRefer to caption
(e) Prot
Refer to captionRefer to caption
(f) Site
Figure 10: Distribution of Rfam clans across task datasets. For each dataset, the top plot excludes unannotated RNAs and the bottom includes them.
Appendix D Timing
We ran our proposed models with a batch size of 8 on 4 Intel i9 CPU cores or on a personal GPU (RTX 3500), and report runtimes and memory requirements in Table 3. Given the modest dataset sizes, runtimes and memory use were not extensively optimized but are unlikely to pose practical limitations. Depending on the task, training a model takes between a few minutes and one hour. Training on a GPU results in a twofold speedup.
Table 3: An overview of the time and memory requirements of training our models for each provided task.
Task Name Time/RNA (CPU ms) Time/RNA (GPU ms) Peak GPU Memory (MB)
RNA_CM 22.4 9.69 60.66
RNA_Go 23.77 11.78 48.84
RNA_IF 10.11 5.49 55.40
RNA_Ligand 21.46 8.65 42.00
RNA_Prot 21.64 9.70 52.04
RNA_Site 23.42 9.95 50.77
RNA_VS 33.55 16.59 41.97
Appendix E Details about representations and models
This section provides a comprehensive overview of the models and RNA representations used in order to foster reproducibility.
E.1 Models
This subsection provides details about the models used.
Bidirectional Long Short-Term Memory (BiLSTM) processes sequential data in both forward and backward directions using two separate Long short-term memory (LSTM) layers. Unlike standard LSTMs that only consider past context, BiLSTM captures both past and future information by concatenating outputs from both directions. This bidirectional approach provides richer contextual representations, making it effective for sequence labeling and classification tasks where the complete sequence is available during processing.
Transformers are attention-based neural networks that process sequential data in parallel rather than sequentially. The core mechanism is self-attention, which allows the model to weigh the importance of different positions in the input sequence when computing representations for each token. Unlike RNNs, transformers can process all sequence positions simultaneously, enabling faster training and better capture of long-range dependencies. The architecture consists of encoder-decoder blocks with multi-head attention, position encodings, and feed-forward layers.
Graph convolutional network (GCN) [Kipf, 2016] is a graph neural network architecture proposing to extend convolution operations to graph data by performing message passing on the 1-hop neighborhood of each node at each layer. The formula of the GCN is the following:
$h^{l+1}=\sigma(Ah^{l}W+b)$
where $h^{l}$ denotes the embedding at the end of layer $l$, $\sigma$ denotes an activation function, $A$ the adjacency matrix of the graph, $W$ the weights of layer $l+1$ and $b$ the bias of layer $l+1$
Relational graph convolutional network (RGCN) [Schlichtkrull et al., 2018] is a refinement of GCN with different edge types, and distinct learnable weight matrices for each edge type. It has the following formula:
$h^{l+1}=\sigma(W_{0}h^{l}+\sum_{r=1}^{R}W_{r}A^{r}h^{l})$
where $R$ is the number of edge types and $A^{r}$ the adjacency matrix relative to edge type $r$
Geometric vector perceptron (GVP) [Jing et al., 2020], initially introduced to learn on protein structures, is an equivariant graph neural network designed to process both scalar and vector node and edge features. In particular, it enables to process 3D coordinates in the learning process. The GVP equations consist in the update equations of both scalar $s^{l}$ and vector $V^{l}$ features at layer $l+1$:
$s^{l+1}=\sigma(W_{m}[\|W_{h}V^{l}\|_{2}||s^{l}])$
$V^{l+1}=\sigma^{+}(\|W_{\mu}W_{h}V^{l}\|_{2})\odot W_{\mu}W_{h}V^{l}$
where $\sigma$ and $\sigma^{+}$ are two activation functions and $W_{h}$ and $W_{\mu}$ are two learnable linear transformations, $\odot$ denotes row-wise multiplication and $[\cdot||\cdot]$ denotes concatenation.
E.2 Representations
We here detail the various representation frameworks proposed in Rnaglib as Representation objects and used in our experiments.
2D representation.
RNA is represented as a graph. Each node corresponds to a nucleotide and each edge encodes either a backbone connection or a canonical base pair. Node features are the one-hot encoding of base identity of nucleotides. Edges are not featurized.
2D+ representation.
RNA is represented as a graph, very similar to the 2D representation. The only difference lies in the presence of three distinct edge types: backbone 3’-5’ connections, backbone 5’-3’ connections and base pairing edges.
2.5D representation.
RNA is represented as a graph whose nodes correspond to a nucleotide and edges encode either a backbone connection or a base pair, be it canonical or non-canonical. This representation is built using structures coming from the Protein Data Bank (PDB). Node features encode the base identity of nucleotides and, optionally, comprise an embedding computed from the RNA language model RNA-FM [Chen et al., 2022]. This graph has distinct edge types. Not only do these edge types distinguish between backbone and base pair connections (as in 2D+ representation), they also account for the geometry of base pair interactions according to Leontis and Westhof categorization.
3D representations.
RNA is represented as a graph which nodes represent its nucleotides. We provide two possibilities for edge definition: either a $k$-nearest neighbor graph (edges are built between a node and its $k$ nearest nodes in the 3D space) or a graph which edges encode either a backbone connection or a canonical base pair (as in 2D and 2D+ representation). In our experiments, we report results for $k$-nearest neighbor graphs with $k=16$. Node scalar features encode the base identity of nucleotides and, optionally, comprise an embedding computed from the RNA language model RNA-FM. Node vector features are unit vectors from nucleotide i-1 to nucleotide i and from nucleotide i to nucleotide i+1) and, optionally, the base pair geometry according to Leontis and Westhof nomenclature (this option is typically used in our GVP-2.5D representation). Edge scalar features include radial basis function encoding of the distance between the two nucleotides forming the edge. As edge vector features, we use the unit vector from nucleotide i to nucleotide j, where i and j are the nucleotides represented by the nodes forming the edge. By default and in our experiment, the 3D coordinates of the nucleotides are assumed to be the 3D coordinates of their phosphate atoms. However, the implementation of the representation is very modular and gives the users the possibility to add custom edge or node features, as well as to use different representative atoms to model nucleotide coordinates.
Appendix F Training details
We provide all the code needed to reproduce our training in the two attached repositories. The code is accompanied by readme documents describing the step-by-step process to reproduce our results. In addition, extensive code documentation is available for all parts of the developed tool. For a simple overview, the most important hyperparameters of the GCNs and RGCNs used in our experiments are summarized in table  4. All our experiments have been performed using a batch size of 8.
Table 4: Hyperparameters used for different RNA tasks in GCN and RGCN models.
Model Representation Layers Hidden Dim Epochs Learning Rate Dropout
RNA_Ligand 2.5D 3 64 20 $1\times 10^{-3}$ 0.5
RNA_CM 2.5D 3 128 40 $1\times 10^{-3}$ 0.5
2D 3 128 40 $1\times 10^{-3}$ 0.5
2D-GCN 2 256 40 $1\times 10^{-3}$ 0.5
RNA_Site 2.5D 4 256 40 $1\times 10^{-3}$ 0.5
2D 2 128 40 $1\times 10^{-4}$ 0.5
2D-GCN 4 256 100 $1\times 10^{-3}$ 0.5
RNA_Prot 2.5D 4 128 40 $1\times 10^{-3}$ 0.2
2D 4 64 40 $1\times 10^{-2}$ 0.2
2D-GCN 5 256 40 $1\times 10^{-2}$ 0.2
RNA_IF 2.5D 3 128 100 $1\times 10^{-4}$ 0.5
RNA_VS 2.5D 3 64/32 100 $1\times 10^{-3}$ 0.2
RNA_GO 2.5D 3 128 20 $1\times 10^{-3}$ 0.5
We report below the hyperparameters used to train GVP models. In the case of GVP models, we don’t tune a unique hidden dimension hyperparameter but four: the number of hidden node scalar features $h_{s}^{N}$, the number of hidden node vector features $h_{V}^{N}$, the number of hidden edge scalar features $h_{s}^{E}$ and the number of hidden edge vector features $h_{V}^{E}$. In the table below, we call Node Dim the $(h_{s}^{N},h_{V}^{N})$ couple and Edge Dim the $(h_{s}^{E},h_{V}^{E})$ couple.
Table 5: Hyperparameters used for different RNA tasks in GVP models.
Model Representation Layers Node Dim Edge Dim Epochs Learning Rate Dropout
RNA_CM GVP 3 (128, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
GVP-2.5D 3 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
RNA_Prot GVP 4 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
GVP-2.5D 4 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
RNA_Site GVP 6 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
GVP-2.5D 6 (32, 2) (32, 1) 40 $1\times 10^{-3}$ 0.5
We report below the hyperparameters used for the sequence models (BiLSTM and Transformer).
Table 6: Hyperparameters used for different RNA tasks in Sequence models.
Model Model Layers Hidden Dim Epochs Learning Rate Dropout
RNA_CM BiLSTM 2 128 500 $1\times 10^{-3}$ 0.5
Transformer 2 32 100 $1\times 10^{-5}$ 0.2
RNA_Prot BiLSTM 2 64 500 $1\times 10^{-3}$ 0.2
Transformer 2 128 100 $1\times 10^{-5}$ 0.2
RNA_Site BiLSTM 2 256 500 $1\times 10^{-3}$ 0.5
Transformer 8 256 100 $1\times 10^{-5}$ 0.2
For all binary tasks, we apply sample weighting to the loss in order to account for the class imbalance. The weights attributed to positive samples are given by the following formula:
$\sqrt{\frac{N_{neg}}{N_{pos}}}$
where $N_{neg}$ is the number of negative samples in the dataset and $N_{pos}$ is the number of positive samples.
For 2D-GCN representations only, we directly use the negative-to-positive ratio as sample weights instead of its square root because we found this reinforcement of the weighting to ease training in this setting.
Appendix G Additional Results
G.1 Additional metrics for the reported models